Ch 13 Ex 13-7 Saved Help Save & Exit Submit Check my work Simon Companys year-end balance sheets follow. Current Yr 1 Yr AgoComplete this question by entering your answers in the tabs below. Required 1A Required 1B Required 2A Required 2B Compute thComplete this question by entering your answers in the tabs below. Required 1A Required 1B Required 2A Required 2B Did the cu(1-a) Compute the current ratio for each of the three years. (1-b) Did the current ratio improve or worsen over the three yea(1-a) Compute the current ratio for each of the three years. (1-b) Did the current ratio improve or worsen over the three yea
Answer #1
Current Ratio
Choose Numerator: / Choose Denominator: = Current Ratio
Current assets / Current liabilities = Current ratio
Current yr. $206,490 / $108,662 = 1.90 to 1
1 Yr. ago $164,112 / $66,173 = 2.48 to 1
2 Yrs. ago $122,276 / $41,805 = 2.92 to 1
1B
Current ratio worsen over three year period.
2A
Acid-Test Ratio
Choose Numerator: / Choose Denominator: = Acid-Test Ratio
Cash + Short-term investments + Current receivables / Current liabilities = Acid-test ratio
Current yr. $26,535 + 0 + $77,669 / $108,662 = 0.96 to 1
1 Yr. ago $31,018 + 0 + $55,355 / $66,173 = 1.31 to 1
2 Yrs. ago $31,037 + 0 + $41,805 / $41,805 = 1.74 to 1
2B
Acid-test ratio worsen over three year period.
Workings:
Current yr. 1 Yr. ago 2 Yrs. ago
Cash 26535 31018 31037
Accounts Receivable 77669 55355 41805
Merchandise Inventory 93825 69597 45880
Prepaid expenses 8461 8142 3554
Current assets 206490 164112 122276